# Optimal Policy Minimum Bayesian Risk

### v1

@@ Abstract
Inference scaling can help LLMs solve complex reasoning problems through extended runtime computation. On top of targeted supervision for long chain-of-thought (long-CoT) generation, purely inference-time techniques such as best-of-N (BoN) sampling, majority voting, or more generally, minimum Bayes risk decoding (MBRD), can further improve LLM accuracy by generating multiple candidate solutions and aggregating over them. These methods typically leverage additional signals in the form of reward models and risk/similarity functions that compare generated samples, e.g., exact match in some normalized space or standard similarity metrics such as Rouge. Here we present a novel method for incorporating reward and risk/similarity signals into MBRD. Based on the concept of optimal policy in KL-controlled reinforcement learning, our framework provides a simple and well-defined mechanism for leveraging such signals, offering several advantages over traditional inference-time methods: higher robustness, improved accuracy, and well-understood asymptotic behavior. In addition, it allows for the development of a sample-efficient variant of MBRD that can adjust the number of samples to generate according to the difficulty of the problem, without relying on majority vote counts. We empirically demonstrate the advantages of our approach on math (MATH-$500$) and coding (HumanEval) tasks using recent open-source models. We also present a comprehensive analysis of its accuracy-compute trade-offs.
@@ 1 Introduction
In this work, we propose an improvement to standard MBRD, termed Optimal Policy MBRD (OP-MBRD), with the following desirable properties:
@@
  • OP-MBRD retains most of the simplicity of BoN and MBRD, while introducing only a single new parameter and remaining compatible with standard MBRD (i.e., beyond the use of simple exact match as in ordinary majority voting).
@@ 2 Related work
Reward-weighted Post-Processing: These can be seen as an enhancement of the previous. They generate $N$ independent sentences in the same way, but utilize a separate model to score the completed outputs e.g. a reward model. Best-of-N (BoN)  (Charniak and Johnson, 2005) is a very common and simple method with proven success to enhance LLM performance  (Yang et al., 2024). These methods also include combinations of majority voting with reward models, which provide improved performance with respect to plain majority voting, e.g., voting verifiers  (Li et al., 2023). These methods can also be expressed as a form of MBRD where pair-wise risk/similarity is enhanced with the output reward. Reward post-processing compensates some of the limitations of pure MV while adding only the overhead of a call to to an external reward model. The method here presented falls into this category. As shown in the next sections it retains the simplicity advantage of similar counterparts such as BoN or voting verifiers, while being derived from well understood principles, providing more robust performance and an efficient version with better performance/compute trade-offs.
@@
Optimal Policy in Reinforcement Learning The method introduced here is also related in its mathematical background to recent works in the reinforcement learning. Methods like proximal policy optimization (PPO)  (Schulman et al., 2017), GDC++  (Korbak et al., 2022) and BRAIn  (Pandey et al., 2024) optimize the KL-controlled reward maximization objective through different approximations. The solution to this problem is the intractable optimal policy, which is the basis of the decoding algorithm presented here. Direct preference optimization (DPO)  (Rafailov et al., 2023) is also derived from the optimal policy and can be proven to be a special case of BRAIn. Rejection sampling from the optimal posterior is also considered in statistical rejection sampling optimization  (Liu et al., 2024) as a method similar to DPO.
@@ 3 Minimum Bayesian Risk Decoding
Here $g()$ is a function that extracts an answer from each model output (which may contain CoT and other tokens), compares them using some normalization, e.g., a symbolic representation, and returns $1$ if they are equal or 00 otherwise. This amounts to selecting the answer that occurs more often in this normalized space. Other forms of MBRD include using symbol-level distances such as Rouge  (Lin, 2004). In some fields like machine translation, evaluation metrics like BLEU  (González-Rubio et al., 2011) or COMET  (Guttmann et al., 2024) are also used. For LLMs, it is straightforward to incorporate a reward model into the risk/similarity computation as
$M(y,y^{\prime},x)=\delta_{g(y,y^{\prime})}\cdot R(y,x),$ (5)
which can substantially improve performance. This can also be viewed as an instance of a voting verifier  (Li et al., 2023). This setup will henceforth be referred to as MBRD (EM*R).
@@ 4.1 Definition
We here propose another way of combining $p(y\mid x)$ and $R(y,x)$ that represents a minimum increase in complexity, while providing interesting properties. Borrowing from Reinforcement Learning, one can define a distribution $q$ that maximizes a expected reward $R(y,x)$ while being close to a reference distribution $p_{R}(y\mid x)$. This can be expressed as the objective
@@ 4.1.1 Computing Expectations with respect to the Optimal Policy
Since softmax is invariant to shifting the logits by a constant, Rao-Blackwellized rejection sampling in Eq. 11 coincides with self-normalized importance sampling (SNIS) (Bengio and Senécal, 2008) with unnormalized weights $p_{n}$. A bias subtracted version of SNIS is also used in BRAIn, but for the purpose of reinforcement learning. We term this last estimator Optimal Policy Minimum Bayesian Risk (OP-MBR) and its maximization OP-MBR Decoding (OP-MBRD).
@@ 4.2 OP-MBRD with a Process Reward Model
where $M$ is the sum of maximum $\tilde{R}$ for each step that ensures we are multiplying probabilities. In practice we normalize the sum of PRM scores by the number of steps $T$. Note that this does not require step-by-step decoding. The outputs are fed to the PRM at the end of generation with appropriate markers i.e. double end of line, and the PRM returns scores for what it considers steps. Another possible interpretation of this formula is ancestral importance sampling of the optimal policy.
@@ 4.3 Efficient OP-MBRD
Since OP-MBRD Rao-Blackwellized rejection sampling and the importance sampling estimators coincide, it may seem that the rejection sampling formulation is redundant. Nevertheless, one can still derive a useful metric from it, the number of expected optimal policy samples for a sample set $\mathcal{S}(N)$
@@
Under the rejection sampling interpretation of our estimator, this gives us a measure of how successful our sampling round was, with a higher $\hat{N}^{\mbox{{\tiny OP}}}$ indicating more samples belong to $p^{*}$. We can use this to derive an efficient version of OP-MBRD, where we fix a desired number of optimal policy samples $N^{\mbox{{\tiny OP}}}$ and we sample repeatedly until $\hat{N}^{\mbox{{\tiny OP}}}\geq N^{\mbox{{\tiny OP}}}$. As it will be shown in the experimental setup, these yields a good prediction of task difficulty for generator-PRM pairs. We will describe those pairs has being well calibrated. We will henceforth refer to this proposed method as OPE-MBRD.
@@ 4.4 Formal Guarantees
Furthermore, we can examine in detail Eqs 35,10 to study what sampling from the optimal policy entails. For cases in which $R=0$ only the log-ratio term remains and OP-MBRD reduces to MBRD from $p_{R}(y\mid x)$, as approximated by self-normalized importance sampling. If we use our generator as reference $p_{R}(y\mid x)=p(y\mid x)$, only the reward term $R$ remains. For a PRM this now will represent MBRD with respect to an energy model proportional to the odds of reaching the correct answer. For an oracle PRM this would assign zero weight to any sample in $S(N)$ not reaching the correct answer, which in the limit guarantees that OP-MBRD would always choose the right answer  [^9] Assuming a perfect PRM, $\lim_{N\to\infty}$ and model assigning non zero probability to the solution.
@@ 5.1 Models and Datasets
To evaluate the methods proposed, we test small and medium LLMs on math and coding tasks. For reproducibility and completeness, we select recent open source models in the 1-20 billion parameter range. For math, we select Alibaba’s Qwen-2.5-math models  (Abdin et al., 2024) sizes 1.5b and 7b as high performing math-specific models. These have a matching process reward model – Qwen-2.5-PRM-7b  (Zhang et al., 2025) – that we also use in our experiments. We also select IBM’s Granite 3.3101010https://huggingface.co/ibm-granite/granite-3.3-8b-instruct models sizes 2b and 8b. These are general models that also exhibit strong math performance compared to, e.g., LLaMa models of the same size (Grattafiori et al., 2024). For the Granite models, we train our own PRM from Granite-3.3 for math. The model was trained with synthetically generated data. The training data consists of step-level correctness annotations, generated using the binary search method of  Luo et al. (2024). The input prompts are sampled from MathInstruct  (Yue et al., 2023), MetaMathQA  (Yu et al., 2023) and NuminaMath  (Li et al., 2024) datasets – the responses are sampled from Granite-3.x, Mixtral-8x22B and Phi4-instruct models. After initial training of the PRM with this data, we use the trained PRM to further filter out low-quality step annotations. We discard samples where step-level correctness annotations do not match the PRM’s assessment of step quality. We then perform a second iteration of PRM training with this higher-quality filtered data. As the upper tier in size we select again Phi-4-instruct (Abdin et al., 2024) (14b) as an additional generator. Finally, we pair Phi4-instruct also with a Phi4-PRM trained the same way. We do not include long-CoT models since we are focusing here on approaches leveraging BoN and MBRD, which leverage independent samples rather than long contexts.
We evaluate all models and methods on MATH-$500$  (Hendrycks et al., 2021; Lightman et al., 2024) and HumanEval  (Chen et al., 2021). MATH-$500$ is a collection of $500$ math competition problems that require detailed step-by-step reasoning to solve. HumanEval consists of $164$ programming problems, each asking to complete a standalone Python function from requirements specified in a docstring. Unit tests are included for each example for automatic evaluation. We use pass@$1$ scores to assess performance on both datasets.
@@ 5.2 Baselines and Methods
As inference scaling baselines we focus on well established single-step algorithms that ensemble multiple samples. We consider BoN using the average PRM score across steps, which was observed to be more performant than other aggregations like minimum in these datasets. We use also two variants of MBRD. First, variants not making use of a PRM or any other parametric scoring function. For MATH-$500$ we use exact match similarity, $M(y,y^{\prime})=\delta_{g(y,y^{\prime})}$. This is often also described as majority voting or consensus and is here referred as MBRD (EM). As text normalizer $g()$ we extract the answer inside the a boxed command, including a prompt to force the model to adhere to such format. For code, exact match performs very poorly, we use Rouge  (Lin, 2004) instead, $M(y,y^{\prime})=\mathrm{rouge}(y,y^{\prime})$. This performed better in our initial test than other alternatives such as Python’s difflib.SequenceMatcher  (Ratcliff and Metzener, 1988; Wei et al., 2025). We refer to this as MBRD (rouge). As parametric MBRD baseline, we used Voting Verifier  (Li et al., 2023), which can be expressed as $M(y,y^{\prime})=\delta_{g(y,y^{\prime})}\cdot\mathrm{mean}\mathrm{PRM}(y,x)$ here referred to as MBRD (EM*R).
@@
As methods proposed, we introduce Optimal Policy variants of the non-parametric MBRD i.e. OP-MBRD (EM) and OP-MBRD (rouge). These use the Rao-Blackwellized rejection sampling (or equivalently importance sampling) to sample from the optimal posterior (see Section 4). All our experiments use $p_{R}(y\mid x)=p(y\mid x)$, which in practice nullifies the effect of the log-ratio term of OP-MBRD. Although the log-ratio term can be proved to equate rejection/importance sampling of $p_{R}$ (see Section 4.4), initial experiments did not show big advantages when using strong teachers for $p_{R}$ and we leave further exploration for future work. This also allowed to estimate the maximum reward in Eq. 35 as the maximum PRM value $1.0$. In addition to the normal variants, we also used the efficient version proposed in Section 4.3, termed OPE-MBRD, which uses the expected amount of accepted samples to decide when to stop sampling. For this we iteratively sampled outputs one by one until a target budget of $N^{\mbox{{\tiny OP}}}=\{1,2,4,8,16,32,64,128,256\}$ optimal samples was met. These experiments are designed to measure the gains in throughput and not in wall-clock time. For the latter, a schedule would have to be designed that uses the observed probability of success to guess a fixed number of samples to be generated next. We leave this for future work.
5.3 Hyperparameters and Variance Reduction
@@
Figure 1: MATH-$500$ test results. Left: pass@1 score as a function of the number of samples per input. Every dot marker is an average over multiple experiment repetitions. A star marker denotes the efficient OP-MBRD represented on the right side. Right: Number of samples OPE-MBRD selects for every example in the test set, sorted from easy to difficult by regular decoding pass@1. Largest standard deviation among all experiments was less than $0.5$.
@@
For hyperparameter tuning, we construct a development set out of NuminaMath [^11] https://huggingface.co/datasets/AI-MO/NuminaMath-CoT. We include a random subset of $500$ question-answer pairs in this set, discarding their CoTs, and making sure they (a) pass simple format check, and (b) are not in MATH-$500$. We set the KL-term weight $\beta$, representing the relative weight of the generator versus the PRM in this set. A value of $\beta=0.1$ was found to be robust across many scenarios and was selected for Qwen and Granite models both for math and code tasks. The only clear exception was Qwen-15b/Qwen-7B-PRM. Results on the dev set indicated that, for this pair, the PRM is much stronger than the generator, and a value of $\beta=0.001$ was selected. Similarly a value of $\beta=0.01$ was selected for Phi-4/Phi-4-PRM. For the OPE-MBRD a maximum number of samples was set as a ×$\times 10$ multiplier of the number of optimal samples selected. For e.g. if we solicited $2$ optimal samples, no more than $20$ real samples would be sampled. This was a simple compromise that helped with badly calibrated generator-PRM pairs, that tend to have spikes in the number of samples solicited. We include the full dev details in Appendix B.
@@
To cover a wide range of inference scaling cases, we produce $N=256$ samples for each input. To reduce variance of results, we always make use of the pool of $256$ samples for all experiments, either for ensembling or experiment repetition. For example, for MBRD (EM) with $16$ samples, we can repeat the experiment $256/16=16$ times. We show the average performance for these $16$ repetitions. Note that for the efficient version of OP-MBRD, OPE-MBRD, the number of samples that constitutes an experiment changes, since the algorithm can select a different number of samples for different generations. We consume blocks of samples of variable size until exhausting the sample pool to construct experiment repetitions. No sample is ever shared across experiments. Standard deviation across all conditions ranged from $0.3$ to $0.5$.
@@
Figure 2: HumanEval test results: Left: pass@1 score as a function of the number of samples per input. Every dot marker is an average over multiple experiment repetitions. A star marker denotes the efficient OP-MBRD represented on the right side. Right: Number of samples OPE-MBRD selects for every example in the test set, sorted from easy to difficult by regular decoding pass@1. Largest standard deviation among all experiments was less than $0.5$.

### v2

@@ Abstract
Inference scaling helps LLMs solve complex reasoning problems through extended runtime computation. On top of long chain-of-thought (long-CoT) models, purely inference-time techniques such as best-of-N (BoN) sampling, majority voting, or more generally, minimum Bayes risk decoding (MBRD), can further improve LLM accuracy by generating multiple candidate solutions and aggregating over them. These methods typically leverage additional signals in the form of reward models and risk/similarity functions that compare generated samples, e.g., exact match in some normalized space or standard similarity metrics such as Rouge. Here we present a novel method for incorporating reward and risk/similarity signals into MBRD. Based on the concept of optimal policy in KL-controlled reinforcement learning, our framework provides a simple and well-defined mechanism for leveraging such signals, offering several advantages over traditional inference-time methods: higher robustness, improved accuracy, and well-understood asymptotic behavior. In addition, it allows for the development of a sample-efficient variant of MBRD that can adjust the number of samples to generate according to the difficulty of the problem, without relying on majority vote counts. We empirically demonstrate the advantages of our approach on math (MATH-$500$) and coding (HumanEval) tasks using recent open-source models. We also present a comprehensive analysis of its accuracy-compute trade-offs.
@@ 1 Introduction
In this work, we propose an enhancement to standard MBRD, termed Optimal Policy MBRD (OP-MBRD), with the following desirable properties:
@@
  • OP-MBRD retains most of the simplicity of BoN and MBRD, while introducing only a single new parameter and remaining compatible with general MBRD (i.e., beyond the use of simple exact match as in ordinary majority voting).
@@ 2 Related work
Reward-weighted Post-Processing: These can be seen as an enhancement of the previous. They generate $N$ independent sentences in the same way, but utilize a separate model to score the completed outputs e.g. a reward model. Best-of-N (BoN)  (Charniak and Johnson, 2005) is a very common and simple method with proven success to enhance performance  (Yang et al., 2024). These methods also include combinations of majority voting with reward models, which provide improved performance with respect to plain majority voting, e.g., voting verifiers  (Li et al., 2023). These methods can also be expressed as a form of MBRD where pair-wise risk/similarity is enhanced with the output reward. Reward post-processing compensates some of the limitations of pure MV while adding only the overhead of a call to to an external reward model. The method here presented falls into this category. As shown in the next sections it retains the simplicity advantage of similar counterparts such as BoN or voting verifiers, while being derived from well understood principles, providing more robust performance and an efficient version with better performance/compute trade-offs.
@@
Optimal Policy Approaches The method introduced here is also related in its mathematical background to recent works using Optimal Policy. Methods like proximal policy optimization (PPO)  (Schulman et al., 2017), Direct preference optimization (DPO)  (Rafailov et al., 2023), GDC++  (Korbak et al., 2022) and BRAIn  (Pandey et al., 2024) utilize the same KL-controlled reward maximization objective but for the purpose of deriving a Reinforcement Learning loss. Rejection Sampling Optimization (RSO) (Liu et al., 2024) modifies DPO to use samples from the optimal posterior via rejection sampling, while Guide Speculative Inference (GSI) Geuter et al. (2025) leverages it for speculative decoding in inference scaling. The former is also a RL learning algorithm while the later is related to our use of log-ratio, see experimenta setup.
@@ 3 Minimum Bayesian Risk Decoding
Here $g()$ is a function that extracts an answer from each model output (which may contain CoT and other tokens), compares them using some normalization, e.g., a symbolic representation, and returns $1$ if they are equal or 0 otherwise. This amounts to selecting the answer that occurs more often in this normalized space. Other forms of MBRD include using symbol-level distances such as Rouge  (Lin, 2004). In some fields like machine translation, evaluation metrics like BLEU  (González-Rubio et al., 2011) or COMET  (Guttmann et al., 2024) are also used. For LLMs, it is straightforward to incorporate a reward model into the risk/similarity computation as
$M(y,y^{\prime},x)=\tilde{M}(y,y^{\prime},x)\cdot R(y,x),$ (5)
where $\tilde{M}$ is the original similarity function. This method can also be viewed as an instance of a voting verifier  (Li et al., 2023). This setup will henceforth be referred to as MBRD (EM*R).
@@ 4.1 Definition
We here propose another way of combining $p(y\mid x)$ and $R(y,x)$ that represents a minimum increase in complexity, while providing interesting properties. Borrowing from Reinforcement Learning, one can define a distribution $q$ that maximizes an expected reward $R(y,x)$ while being close to a reference distribution $p_{R}(y\mid x)$. This can be expressed as the objective
@@ 4.1.1 Computing Expectations with respect to the Optimal Policy
Since softmax is invariant to shifting the logits by a constant, Rao-Blackwellized rejection sampling in Eq. 11 coincides with self-normalized importance sampling (SNIS) (Bengio and Senécal, 2008) with unnormalized weights $p_{n}$. We term this last estimator Optimal Policy Minimum Bayesian Risk (OP-MBR) and its maximization OP-MBR Decoding (OP-MBRD). A full derivation can be found in Appendix E.
@@ 4.2 OP-MBRD with a Process Reward Model
where $M$ is the sum of maximum $\tilde{R}$ for each step. In practice we normalize the sum of PRM scores by the number of steps $T$. Note that this does not require step-by-step decoding. The outputs are fed to the PRM at the end of generation with appropriate markers i.e. double end of line, and the PRM returns scores for what it considers steps. Another possible interpretation of this formula is ancestral importance sampling of the optimal policy.
@@ 4.3 Efficient OP-MBRD
Since OP-MBRD Rao-Blackwellized rejection sampling and importance sampling estimators coincide, it may seem that the rejection sampling formulation is redundant. Nevertheless, one can still derive a useful metric from it, the number of expected optimal policy samples for a sample set $\mathcal{S}(N)$
@@
Under the rejection sampling interpretation of our estimator, this gives us a measure of how successful our sampling round was, with a higher $\hat{N}^{\mbox{{\tiny OP}}}$ indicating more samples belong to $p^{*}$. We can use this to derive an efficient version of OP-MBRD, where we fix a desired number of optimal policy samples $N^{\mbox{{\tiny OP}}}$ and we sample repeatedly until $\hat{N}^{\mbox{{\tiny OP}}}\geq N^{\mbox{{\tiny OP}}}$. As it will be shown in the experimental setup, these yields a good prediction of task difficulty for generator-PRM pairs. We will describe those pairs has being well calibrated and henceforth refer to this proposed method as OPE-MBRD.
@@ 4.4 Formal Guarantees
Furthermore, we can examine in detail Eqs 35,10 to study what sampling from the optimal policy entails. For cases in which $R=0$ only the log-ratio term remains and OP-MBRD reduces to MBRD from $p_{R}(y\mid x)$, as approximated by self-normalized importance sampling. If we use our generator as reference $p_{R}(y\mid x)=p(y\mid x)$, only the reward term $R$ remains. For a PRM this now will represent MBRD with respect to an energy model proportional to the odds of reaching the correct answer. For an oracle PRM this would assign zero weight to any sample in $S(N)$ not reaching the correct answer, which in the limit guarantees that OP-MBRD always chooses the right answer  [^9] Assuming a perfect PRM, $\lim_{N\to\infty}$ and model assigning non zero probability to the solution.
@@ 5.1 Models and Datasets
To evaluate the methods proposed, we test small and medium LLMs on math and coding tasks. For reproducibility and completeness, we select recent open source models in the 1-20 billion parameter range. For math, we select Alibaba’s Qwen-2.5-math Instruct models  (Abdin et al., 2024) sizes 1.5b and 7b as high performing math-specific models. These have a matching process reward model – Qwen-2.5-PRM-7b  (Zhang et al., 2025) – that we also use in our experiments. We also select IBM’s Granite 3.3101010https://huggingface.co/ibm-granite/granite-3.3-8b-instruct models sizes 2b and 8b. These are general models that also exhibit strong math performance compared to, e.g., LLaMa models of the same size (Grattafiori et al., 2024). For the Granite models, we train our own PRM from Granite-3.3 for math. The model was trained with synthetically generated data. The training data consists of step-level correctness annotations, generated using the binary search method of  Luo et al. (2024). The input prompts are sampled from MathInstruct  (Yue et al., 2023), MetaMathQA  (Yu et al., 2023) and NuminaMath  (Li et al., 2024) datasets – the responses are sampled from Granite-3.x, Mixtral-8x22B and Phi4-instruct models. After initial training of the PRM with this data, we use the trained PRM to further filter out low-quality step annotations. We discard samples where step-level correctness annotations do not match the PRM’s assessment of step quality. We then perform a second iteration of PRM training with this higher-quality filtered data. As the upper tier in size we select again Phi-4-instruct (Abdin et al., 2024) (14b) as an additional generator. Finally, we pair Phi4-instruct also with a Phi4-PRM trained the same way. We do not include long-CoT models since we are focusing here on approaches leveraging BoN and MBRD, which leverage independent samples rather than long contexts.
We evaluate all models and methods on MATH-$500$  (Hendrycks et al., 2021; Lightman et al., 2024) and HumanEval  (Chen et al., 2021). MATH-$500$ is a collection of $500$ math competition problems requiring step-by-step reasoning to solve. HumanEval consists of $164$ programming problems, each asking to complete a standalone Python function from its docstring. Unit tests are included for each example for automatic evaluation. We use pass@$1$ scores to assess performance on both datasets.
@@ 5.2 Baselines and Methods
As inference scaling baselines we focus on well established single-step algorithms. We consider BoN using the average PRM score across steps, which was observed to be more performant than other aggregations like minimum in these datasets. We use also two variants of MBRD. First, variants not making use of a PRM or any other parametric scoring function. For MATH-$500$ we use exact match similarity, $M(y,y^{\prime})=\delta_{g(y,y^{\prime})}$. This is often also described as majority voting and is here termed MBRD (EM). As text normalizer $g()$ we extract the answer inside the boxed command. For code, exact match performs very poorly, we use RougeL  (Lin, 2004) instead, $M(y,y^{\prime})=\mathrm{rougeL}(y,y^{\prime})$, which performed best in initial tests. We refer to this as MBRD (rouge). As parametric MBRD baseline, we used Voting Verifier  (Li et al., 2023), which can be expressed as $M(y,y^{\prime})=\delta_{g(y,y^{\prime})}\cdot\mathrm{mean}\mathrm{PRM}(y,x)$ here termed MBRD (EM*R).
@@
As methods proposed, we introduce Optimal Policy variants of the non-parametric MBRD i.e. OP-MBRD (EM) and OP-MBRD (rouge). These use the Rao-Blackwellized rejection sampling (or equivalently importance sampling) to sample from the optimal posterior (see Section 4). All our experiments use $p_{R}(y\mid x)=p(y\mid x)$, which in practice nullifies the effect of the log-ratio term of OP-MBRD. Although the log-ratio term can be proved to equate rejection/importance sampling of $p_{R}$ (see Section 4.4), initial experiments show no advantage when using strong teachers for $p_{R}$. We provide an analysis in Appendix C. This also allowed to estimate the maximum reward in Eq. 35 as the maximum PRM value $1.0$. In addition to the normal variants, we also used the efficient version proposed in Section 4.3, termed OPE-MBRD, which uses the expected amount of accepted samples to decide when to stop sampling. For this we iteratively sampled outputs one by one until a target budget of $N^{\mbox{{\tiny OP}}}=\{1,2,4,8,16,32,64,128,256\}$ optimal samples was met. These experiments are designed to measure the gains in throughput and not in wall-clock time. For the latter, a schedule would have to be designed that uses the observed probability of success to guess a fixed number of samples to be generated next. We leave this for future work.
5.3 Hyperparameters and Experiment Repetition
@@
Figure 1: MATH-$500$ test results. Left: pass@1 score as a function of the number of samples per input with average and standard deviation displayed. A dashed line denotes the efficient version (OPE-MBRD). A star marker denotes the OPE-MBRD configuration represented on the right side. Right: Number of samples OPE-MBRD selects for every example in the test set, sorted from easy to difficult by regular decoding pass@1.
@@
For hyperparameter tuning, we construct a development set out of NuminaMath [^11] https://huggingface.co/datasets/AI-MO/NuminaMath-CoT. We include a random subset of $500$ question-answer pairs in this set, discarding their CoTs, and making sure they (a) pass simple format check, and (b) are not in MATH-$500$. We set the KL-term weight $\beta$, representing the relative weight of the generator versus the PRM in this set. A value of $\beta=0.1$ was found to be robust across many scenarios and was selected for Qwen, Granite and Phi-4 models both for math and code tasks. The only clear exception was Qwen-1.5b/Qwen-7B-PRM. Results on the dev set indicated that, for this pair, the PRM is much stronger than the generator, and a value of $\beta=0.001$ was selected. For the OPE-MBRD a maximum number of samples was set as a $\times 10$ multiplier of the number of optimal samples selected. For example, if we solicited $2$ optimal samples, no more than $20$ real samples would be sampled. This was a simple compromise that helped with badly calibrated generator-PRM pairs, that tend to have spikes in the number of samples solicited. We include the full dev details in Appendix B.
@@
We investigate inference scaling up to $N=256$ samples per input. To reduce variance of results, we always make use of the pool of $256$ samples for all experiments, either for ensembling or experiment repetition.  [^12] For example, for MBRD (EM) with $N=16$ samples, we can repeat the experiment and average scores over $256/16=16$ repetitions. Note that for the efficient version of OP-MBRD, OPE-MBRD, the number of samples that constitutes an experiment changes, since the algorithm can select a different number of samples for different generations. For this, we consume blocks of samples of variable size until exhausting the sample pool to construct experiment repetitions. No sample is ever shared across experiments. To obtain a final average and standard deviation, we further repeat the experiments above $3$ times.
@@
Figure 2: HumanEval test results: Left: pass@1 score as a function of the number of samples per input with average and standard deviation displayed. A dashed line denotes the efficient version (OPE-MBRD). A star marker denotes the OPE-MBRD configuration represented on the right side. Right: Number of samples OPE-MBRD selects for every example in the test set, sorted from easy to difficult by regular decoding pass@1.